• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuss

    Log entry time 00:23:05 on March31,2000

    Entry number 37356

    keyword=Shift summary

    Thu March 30 evening-shift log:

    target change from 5K to 15K coolant was done at 17:00. Some problems with the HRSE NMR regulation was fixed by Javier.
    Called for sweep at about 17:40.
    Beam was available at 19:05.
    Performed HARP scan to tune the beam spot size, final result 105µmx98µm.
    Started boiling studies. Terminated, because elastic peak is off center in the +5% kinematics. OTR done, width 290µm.
    Start with the delta scan